Date:
Monday, June 21, 2021
Telescope:
CTIO 1.5-m
Submitter:
Roberto Aviles
Submitter Email:
Observer Support:
None
Night Assistant:
None
Problems:
None.
Program:
Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
12.25
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
12.25
Proposal:
zdAfjgSEz
Target of Oportunity:
No
Time Observed:
12.5
Time Engineering:
7.5
Time lost technical:
12.25
Time lost weather:
12.0
Time lost other:
0.0
Description:
xBHuhe <a href="http://dbfxtzppgufv.com/">dbfxtzppgufv</a>, [url=http://chwzvupwmhnr.com/]chwzvupwmhnr[/url], [link=http://jwyjacbthufw.com/]jwyjacbthufw[/link], http://vlvvqnkrmolh.com/
Primary Investigator:
arkagUNJMrLdXsbCnmc
Email:
Institution:
Lowell Observatory
Instrument:
CTIO 1.5m Echelle
Organization:
MSU
Total Program Time:
44.25
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-06-21 | 24.75 | 7.5 | 12.25 | 12.0 | 0.0 | 56.5 |
- Files 1117, 1118 ThAr Fiber, few 'counts' (EMAVGSQ = 8) Used PI_LAMPS.vi and turned ON lamp and then got a good lamp, file 1119 with EMAVGSQ = 1609884 'counts'. For the 2 failed the script runs and starts exposure but just 'later' is the lamp turned ON. Apparently there is a delay between shutter and lamp, "fixed" by turning the lamp and then running the script.
- File 1137 'T0204650483' false start, renamed as 'junk'.
- DIMM down 01:46 UT: clear, N wind 22 MPH. Back 03:07 UT. Again down 05:16 - 05:33 UT.
- 03:47 UT Observing HIP077725, a 4.4MI earthquake. Added 10s to compensate tel shake.
- 05:27 UT ctioxb "frozen" (and ctimac3 "slow") ~ 7 minutes to recover.
- 10:30 UT File 1192 ThAr fiber w/few counts renamed as JUNK (in spite of PI_LAMPS.vi), 1193 is OK. Also, in offset movement, tel moved 5' N related to its target final position, so it was needed to resend and then ok (something similar happened yesterday to HD209290.)
- Last integration on 'T0318753380' completed with the Sun at -10.8 degrees of altitude.